# Generalized Spectral Clustering of Low-Inertia Power Networks

### v1

@@ Appendix B Robustness of the Small Eigenvalues of Linearized System Dynamics Matrix
Notice that the pair $(L,D)$ is definite for any $L$ when the matrix $D$ is positive definite.
We consider symmetric perturbations $\Delta L$ and $\Delta D$ to these matrices, and define the perturbed matrices as $\tilde{L}=L+\Delta L$ and $\tilde{D}=D+\Delta D$, let $\sigma(\tilde{L},\tilde{D})=\{\langle\tilde{\alpha}_{1},\tilde{\beta}_{1}\rangle,\langle\tilde{\alpha}_{2},\tilde{\beta}_{2}\rangle,\ldots,\langle\tilde{\alpha}_{n},\tilde{\beta}_{n}\rangle\}$ denote it’s eigenvalues. The chordal distance between the non-zero generalized eigenvalues $\langle\alpha,\beta\rangle$ and $\langle\tilde{\alpha},\tilde{\beta}\rangle$ is given by
$\text{dist}_{c}(\langle\alpha,\beta\rangle,\langle\tilde{\alpha},\tilde{\beta}\rangle)=\frac{|\tilde{\beta}\alpha-\tilde{\alpha}\beta|}{\sqrt{|\alpha|^{2}+|\beta|^{2}}\sqrt{|\tilde{\alpha}|^{2}+|\tilde{\beta}|^{2}}}.$
We denote the stacked perturbation matrix by $\Delta L,\Delta D]\in\mathbb{R}^{2n\times n}$, and use $\Delta L,\Delta D]||_{2}$ to denote its spectral norm.
@@
If we assume that the damping coefficients of the oscillators are constant and consider a perturbation to the laplacian of the dynamic graph $L$, this corresponds to a linearization of the system dynamics about a different equilibrium near the nominal operating point,
@@
[…]
where $\rho(\Delta L)=\max_{i}|\lambda_{i}(\Delta L)|$ is the spectral radius of $\Delta L$.
[…]
$\max_{1\leq j\leq n}\text{dist}_{c}(\langle\alpha_{j},\beta_{j}\rangle,\langle\tilde{\alpha}_{\tau(j)},\tilde{\beta}_{\tau(j)}\rangle)$ $\leq\frac{\rho(\Delta L)}{D_{\min}}.$
If we consider the perturbation $\Delta L$ to represent a change in the laplacian of the dynamic graph at an operating point (different from the nominal), then the worst-case change in the eigenvalues of the linearized dynamics equation of the coupled oscillator system is upper bounded by the ratio of the spectral radius of the perturbation matrix $\Delta L$ to the minimum damping coefficient in the oscillators.
An implication of this inequality for the number of clusters in the network is that, provided the pair $(L,D)$ is not ill-conditioned (that is $\mu(L,D)$ is not too small), if the ratio $\rho(\Delta L)/D_{\min}$ is smaller than the chordal distance between the $k$th and (k+1)(k+1)st eigenvalues, then the linearized dynamics matrix at the new operating point with the laplacian of the dynamic graph given by $L+\Delta L$ also has $k$ small eigenvalues – indicating the presence of $k$ clusters in the network.
Appendix C Perturbation bound on the Eigenspace
@@
$\begin{bmatrix}Z_{1}^{H}\\ Z_{2}^{H}\end{bmatrix}L[Z_{1},Z_{2}]=\begin{bmatrix}L_{1}&0\\ 0&L_{2}\end{bmatrix}$
@@
$\begin{bmatrix}Z_{1}^{H}\\ Z_{2}^{H}\end{bmatrix}D[Z_{1},Z_{2}]=\begin{bmatrix}D_{1}&0\\ 0&D_{2}\end{bmatrix}$
@@
$\delta=\min_{i,j}\{\text{dist}_{c}(\langle\alpha_{i},\beta_{i}\rangle,\langle\tilde{\alpha}_{j},\tilde{\beta}_{j}\rangle):\;$ $\langle\alpha_{i},\beta_{i}\rangle\in\sigma(L_{1},D_{1}),$
$\langle\tilde{\alpha}_{j},\tilde{\beta}_{j}\rangle\in\sigma(\tilde{L}_{2},\tilde{D}_{2})\}.$
If $\delta>0$, then
@@
and for any unitary-invariant norm, $\sin\Theta_{1}||$ is a measure of the difference between the subspaces $\mathcal{R}(Z_{1})$ and the $\mathcal{R}(\tilde{Z}_{1})$.
Given an $\ell$-dimensional eigenspace $\mathcal{R}(Z_{1})$ of the pair $(L,D)$ if the corresponding generalized eigenvalues of $(L_{1},D_{1})$ are well separated from the generalized eigenvalues of $(\tilde{L}_{2},\tilde{D}_{2})$, where the separation is measured by $\delta$, then the difference between the eigenspaces $\mathcal{R}(Z_{1})$ and $\mathcal{R}(\tilde{Z}_{1})$ is upper bounded the inequality (7). Moreover, for any simple eigenspace of $(L,D)$, i.e. one dimensional eigenspace, the more separated the eigenvalue corresponding to this $1$-dimensional eigenspace is from n1n-1 eigenvalues of $(\tilde{L}_{2},\tilde{D}_{2})$, i.e. a large value of $\delta$ and a small value of $1/\delta$ the more stable the eigenspace $\mathcal{R}(Z_{1})$ is to perturbations in $L$ and $D$. Similarly, the subspace spanned by the first $k$ eigenvectors of $(L,D)$ is stable to perturbations $(\Delta L,\Delta D)$, if the $k$ eigenvalues of $(L_{1},D_{1})$ are well separated from nkn-k eigenvalues of $(\tilde{L}_{2},\tilde{D}_{2})$.
Appendix D Coherence of Oscillators
We define $e_{i}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$ to be the deviation of the angular frequency of the $i$th oscillator following a disturbance from the synchronous frequency $\omega_{\text{sync}}$ at time $t$ as
$e_{i}(t)=\dot{\delta}_{i}(t)-\omega_{\text{sync}}$
where $\omega_{\text{sync}}=(\sum_{i}\omega_{i})/(\sum_{i}D_{i})$, $\omega_{i}$ is the natural frequency and $D_{i}$ time constant of the $i$th oscillator.
The $\mathcal{L}_{2}$-norm of the function $e_{i}(t)$ over the finite interval $t_{0},t_{f}]$ is defined as
$e_{i}(t)||=\left(\int_{t_{0}}^{t_{f}}e_{i}(t)^{2}dt\right)^{1/2}.$
And for $i\neq j$, the inner product between the functions $e_{i}(t)$ and $e_{j}(t)$ is defined as
$\langle e_{i}(t)|e_{j}(t)\rangle=\int_{t_{0}}^{t_{f}}e_{i}(t)e_{j}(t)dt.$
The coherency between oscillators $i$ and $j$ can be quantified using their normalized inner product as) as the cosine of the angle between the two frequencies denoted by
[…]
Notice that $\cos(\theta_{ij})\in[-1,1]$, with perfectly coherent pairs of oscillators having coherence values $\cos(\theta_{ij})\rightarrow 1$. The coherence value between $\delta_{i}$ and $\delta_{j}$ does not depend on the magnitude of the deviations $\omega_{\text{sync}}$, just the the alignment of the functions.

### v2

@@ Appendix A A Perturbation Bound on Eigenvalues
Consider symmetric perturbations $\Delta L$ and $\Delta D$ to these matrices. Define $\widehat{L}=L+\Delta L$ and $\widehat{D}=D+\Delta D$, and let $\sigma(\widehat{L},\widehat{D})=\{\widehat{\lambda}_{1},\widehat{\lambda}_{2},\ldots,\widehat{\lambda}_{n}\}$ denote it’s eigenvalues. The chordal distance between the non-zero generalized eigenvalues $\lambda$ and $\widehat{\lambda}$ is defined as
$\text{dist}_{c}(\lambda,\widehat{\lambda})=\frac{|\lambda-\widehat{\lambda}|}{\sqrt{|\lambda|^{2}+1}\sqrt{|\widehat{\lambda}|^{2}+1}}.$
@@
where $\Delta L,\Delta D]||_{2}$ is the operator $2$-norm of the stacked perturbation matrix. If we assume that the damping coefficients of the oscillators are constant, then linearizing the network around different operating points correspond to symmetric perturbations to the Laplacian matrix of the dynamic graph. It follows that the worst-case change in the eigenvalues of the pair $(L,D)$ (and the eigenvalues of the system dynamics matrix $J$) is bounded as
@@
where $\rho(\Delta L)=\max_{i}|\lambda_{i}(\Delta L)|$ is the spectral radius of the perturbation $\Delta L$. Therefore, provided the pair $(L,D)$ is definite, the change in the eigenvalues is bounded by the ratio $\rho(\Delta L)/\min_{i}D_{ii}$.
Appendix B Perturbation bound on the Eigenspace
@@
$\begin{bmatrix}Z_{1}^{H}\\ Z_{2}^{H}\end{bmatrix}L[Z_{1},Z_{2}]=\begin{bmatrix}L_{1}&\mathbf{0}\\ \mathbf{0}&L_{2}\end{bmatrix}$
@@
$\begin{bmatrix}Z_{1}^{H}\\ Z_{2}^{H}\end{bmatrix}D[Z_{1},Z_{2}]=\begin{bmatrix}D_{1}&\mathbf{0}\\ \mathbf{0}&D_{2}\end{bmatrix}$
@@
$\delta=\min_{i,j}\{\text{dist}_{c}(\lambda_{i},\widehat{\lambda}_{j}):\;$ $\lambda_{i}\in\sigma(L_{1},D_{1}),\;\widehat{\lambda}_{j}\in\sigma(\widehat{L}_{2},\widehat{D}_{2})\}.$
If $\delta>0$, then the difference between the subspaces $\mathcal{R}(Z_{1})$ and $\mathcal{R}(\widehat{Z}_{1})$
@@
Given an $\ell$-dimensional eigenspace $\mathcal{R}(Z_{1})$ of the pair $(L,D)$ if the corresponding eigenvalues of $(L_{1},D_{1})$ are well separated from the eigenvalues of $(\widehat{L}_{2},\widehat{D}_{2})$, where the separation is measured by $\delta$, then the difference between the eigenspaces $\mathcal{R}(Z_{1})$ and $\mathcal{R}(\widehat{Z}_{1})$ is bounded as in (8). Moreover, for any simple eigenspace of $(L,D)$, the more separated the eigenvalue corresponding to this $1$-dimensional eigenspace is from n1n-1 eigenvalues of $(\widehat{L}_{2},\widehat{D}_{2})$ as measured by $\delta$ the more stable the eigenspace $\mathcal{R}(Z_{1})$ is to perturbations in $L$ and $D$. Similarly, the subspace spanned by the first $k$ eigenvectors of $(L,D)$ is stable to perturbations, if the corresponding $k$ eigenvalues of $(L_{1},D_{1})$ are well separated from nkn-k eigenvalues of $(\widehat{L}_{2},\widehat{D}_{2})$.
Acknowledgment
The authors would like to thank Prof. David Bindel for the helpful conversation on Matrix Perturbation Theory during the finalization of this work.